You are going to value Lauryn’s Doll Co. using the FCF model. After consulting various sources, you find that Lauryn's has a reported equity beta of 1.5, a debt-to-equity ratio of 0.6, and a tax rate of 30 percent. Assume a risk-free rate of 5 percent and a market risk premium of 9 percent. Lauryn’s Doll Co. had EBIT last year of $48 million, which is net of a depreciation expense of $4.8 million. In addition, Lauryn's made $5.5 million in capital expenditures and increased net working capital by $1.8 million. Assume the FCF is expected to grow at a rate of 3 percent into perpetuity. What is the value of the firm? (Do not round intermediate calculations. Enter your answer in millions rounded to 2 decimal places.)
Answer #1

Using CAPM Model,

Required Rate = Rf + Beta(Rm - Rf)

Required Rate = 0.05 + 1.50(0.09)

Required Rate = 18.50%

Calculating FCF,

FCF0 = 48(1 - 0.30) + 4.8 - 5.5 - 1.8

FCF0 = $31.10 million

So,

Firm Value = FCF0(1 + g)/(r - g)

Firm Value = 31.10(1.03)/(0.185 - 0.03)

Firm Value = $206.44 million